OpenStudy (anonymous):

-3 + 8x - 5 = -8 solve for x

OpenStudy (haseeb96):

x=0 is the correct answer -3 +8x -5 =-8 8x -8 =-8 8x=-8 +8 8x =0 x=0

OpenStudy (anonymous):

No solution, you can't divide.